### v1

# Comment on ”Monochromatization interaction region optics design for direct s-channel Higgs production at FCC-ee”

Abstract
The original article [1] can be logically divided into two parts: 1) the selection of main parameters for monochromatization and 2) interaction region optics design; the comment pertains only to the first part. The authors of [1] state that ”The purpose of this paper is to report on the development of realistic IR optics designs for monochromatization at the FCC-ee”. However, the proposed parameters do not seem very realistic and raise many questions.
keywords:
FCC-ee , Monochromatization , Luminosity , Centre-of-mass energy spread , Beamstrahlung
journal: Nuclear Instruments and Methods in Physics Research A
[…]
organization=Northern Illinois University, addressline=1425 W. Lincoln Hwy., city=DeKalb, postcode=60115, state=IL, country=USA
1 Piwinski angle and collision overlap area
The Piwinski angle is not actually an angle, but a dimensionless parameter. Therefore, the indication that it is measured in radians (Tables 1 and 2) is inappropriate.
In general, the collision overlap area [3] (that is Ref. [14] in [1]) is expressed as $L_{i}=\sigma_{z}\big/\sqrt{1+\varphi^{2}}$. As the authors rightly noted, the $L_{i}\big/\beta_{y}^{*}$ ratio characterizes the hourglass effect. But only for $\varphi\gg 1$ the simplification of the form $L_{i}\approx 2\sigma_{x}^{*}\big/\theta_{c}$ is valid. However, the typical value of Piwinski angle in Tables 1 and 2 (with monochromatization) is $\varphi\approx 1$, so the exact expression for $L_{i}$ must be used. Accordingly, the line ”Overlap collision …” in the Tables should be corrected, which will also affect $R_{hg}$ and the calculated luminosity.
2 Synchrotron tune and $t\bar{t}$ optics
The value and success of the direct s-channel Higgs production experiment depend significantly on the accuracy of the centre-of-mass energy determination, which requires precise beam energy calibration. The most accurate method that everyone is counting on is based on measuring the spin precession frequency. These measurements are performed using resonant depolarization, which in turn imposes stringent requirements on the synchrotron tune.
The modulation of the spin precession frequency by synchrotron oscillations [4] is characterized by the synchrotron modulation index: $\zeta=\nu_{0}\sigma_{\delta}\big/Q_{s}$, where $\nu_{0}\approx 142$ is the spin tune at 62.5 GeV and $\sigma_{\delta}\approx 5.5\cdot 10^{-4}$. For a reliable determination of the spin frequency, one should aim for $\zeta<1.5$, and in any case $\zeta<2$ must be satisfied. It follows that $Q_{s}$ should be at least greater than 0.04, and preferably greater than 0.05. In principle, this is possible in $t\bar{t}$ optics, but it will require raising the RF voltage from 0.17 to 1.01.21.0-1.2 GV. As a consequence, $\sigma_{z}$ will decrease several times, greatly increasing beamstrahlung and making this option less attractive. In any case, many of the values in Table 2 should change significantly.
3 Vertical dispersion
The time available for direct s-channel Higgs production experiment will be limited. Therefore, it is critical to have a sufficiently high luminosity to obtain the required number of events. The dependence of the Higgs event rate on luminosity and $\sigma_{W}$ can be expressed as $\dot{N}_{H}\propto L\Big/\sqrt{\Gamma_{H}^{2}+\sigma_{W}^{2}}$, where $\Gamma_{H}\approx 4.1$ MeV. The original idea was to increase $\dot{N}_{H}$ by monochromatization in the hope that the reduction in the luminosity would be small. Then the goals became more realistic: $L$ can decrease in approximately the same proportion as $\sigma_{W}$, so that $\dot{N}_{H}$ remains almost unchanged. But if monochromatization leads to a noticeable (several times) decrease in $\dot{N}_{H}$, then it loses its meaning.
With this in mind, let us consider the scheme with $D_{y}^{*}\neq 0$ again. On the one hand, it is not hindered by the horizontal crossing angle, and this is the main advantage. On the other hand, increasing $\sigma_{y}^{*}$ leads to a decrease in $L$ by a factor of $\lambda$ (see Eq. (2) in [1]), which is still within acceptable limits. But this works only without BS, that is, for a small bunch population and, consequently, at low luminosity. And taking BS into account, $\dot{N}_{H}$ inevitably falls several times. The advantage of horizontal dispersion over vertical one becomes clear if we take into account the fact that for flat beams, increasing $\sigma_{x}^{*}$ weakens BS, while increasing $\sigma_{y}^{*}$ does not. That is why the option with vertical dispersion was not even considered in the CDR [5].
The results presented in [1] also demonstrate a multiple reduction in $\dot{N}_{H}$ in the case of $D_{y}^{*}\neq 0$. Therefore, to avoid misleading readers, it would be useful to make a clear and distinct conclusion that such a scheme should be rejected.
4 Monochromatization factor
Eq. (3) in [1] is obviously incorrect. To understand this, it suffices to consider the limiting case $D_{y}^{*}=0$, $\varphi\gg 1$ and $D_{x}^{*}\sigma_{\delta}\big/\sqrt{\varepsilon_{x}\beta_{x}^{*}}\gg\varphi$. Then we get $\lambda\gg 1$, which cannot be the case for $\varphi\gg 1$ – this is clear from the geometry shown in Fig. 2 in [1]. Everything is explained simply: the authors took Eq. (3) from [2] (that is Ref. [13] in [1], see Eq. (17) there), but did not take into account that the Piwinski angle in [2] is defined incorrectly: instead of the full horizontal beam size $\sigma_{x}^{*}$, the betatron size $\sigma_{x\beta}^{*}=\sqrt{\varepsilon_{x}\beta_{x}^{*}}$ was substituted into the formula for determining $\varphi$; this issue is considered in detail in [3].
To correct Eq. (3), it is necessary to multiply $\varphi^{2}$ by the factor $(D_{x}^{*2}\sigma_{\delta}^{2}\big/(\varepsilon_{x}\beta_{x}^{*})+1)$. From this, in particular, it follows that $\sigma_{W}$ should increase from 26.4 to about 37.9 MeV for ZH4IP and from 20.2 to about 41.8 MeV for ZH2IP, making the latter scheme less attractive. These values do not agree well with the simulation results presented in Table 3, which raises questions about this simulation.
5 Vertical emittance
In Section 3.3, the authors reasonably noted that $\varepsilon_{y}$ was calculated with the assumed FCC-ee coupling ratio $\varepsilon_{y}\big/\varepsilon_{x}=0.2\%$. If $D_{x}^{*}\neq 0$, $\varepsilon_{x}$ will increase several times due to beamstrahlung, and it would be too optimistic to assume that $\varepsilon_{y}$ will remain unchanged. For example, $\varepsilon_{x}$ increased about 3.7 times for ZH4IP. Under the conservative assumption that about half of this increase is converted to $\varepsilon_{y}$ with the same $0.2\%$ ratio, the luminosity will decrease by approximately 1.5 times.
6 Parameter optimization
The Abstract of [1] contains a phrase that is perceived as a promise that will be fulfilled: ”This measurement is significantly facilitated if the CM energy spread of $e^{+}e^{-}$ collisions can be reduced to a level comparable to the natural width of the Higgs boson, $\Gamma_{H}$ = 4.1 MeV, without substantial loss in luminosity”. And here is the result: $\sigma_{W}$ became $\sim$38 MeV, while the luminosity decreases by 2.3 times (taking into account the aforementioned factor of 1.5). Is this what we were aiming for?
Obviously, to get $\lambda\gg 1$ in a collision scheme with crossing angle and without crab crossing, we need to make $\varphi\ll 1$. To achieve this, $\sigma_{z}$ can be reduced (for example, $\sim$1.5 times) by increasing the RF voltage and $D_{x}^{*}$ should be significantly increased. The statement about a ”minimum bunch spacing of 25 ns” also raises questions. Perhaps this limitation exists at low energy (Z-pole, 45.6 GeV), where the total beam current is 3.5 times higher. However, at 62.5 GeV things become easier, so decreasing $N_{b}$ (along with increasing the number of bunches $n_{b}$) can be a useful option.
An adequate example of parameter optimization can be found in the CDR [5], where the emphasis was on achieving the required $\sigma_{W}$ and luminosity values. Of course, much has changed since then, including the perimeter, the number of IPs, the arc cell type at low energies and hence $\alpha_{C}$, $\varepsilon_{x}$, etc. So, it will be necessary to re-optimize everything, and then, what will be the most difficult part, try to implement it in the IR optics design.
Declaration of competing interest
The author declares that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This manuscript has been authored by FermiForward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.